Yuki (age 45 at year-end) has been contributing to a traditional IRA for years (all deductible contributions), and her IRA is now worth $46,000. She is trying to decide whether she should roll over her traditional IRA into a Roth IRA. Her current marginal tax rate is 25 percent. She plans to withdraw the entire balance of the account in 20 years and she expects to earn a before-tax rate of return of 5.6 percent on her retirement accounts and a 4.6 percent after-tax rate of return on all investments outside of her retirement accounts. For each of the following alternative scenarios, indicate how much more or less Yuki will accumulate after taxes in 20 years if she rolls over her traditional IRA into a Roth IRA. Be sure to include the opportunity cost of having to pay taxes on the rollover. (Round Future value factor to 4 decimal places. Round final answers to the nearest whole dollar amount.) a.When she withdraws the retirement funds in 20 years, she expects her marginal tax rate to be 35 percent. Accumulation if she keeps funds in traditional IRA Accumulation if she rolls funds into Roth IRA accumulation if she rolls over traditional IRA into Roth IRA b.When she withdraws the retirement funds in 20 years, she expects her marginal tax rate to be 20 percent. Accumulation if she keeps funds in traditional IRA Accumulation if she rolls funds into Roth IRA accumulation if she rolls over traditional IRA into Roth IRA
Answer #1

a.

Accumulation if she keeps funds in traditional IRA

$88910 [$46,000 × 1.05620 × (1 - .35)]

Accumulation if she rolls funds into Roth IRA:     $108514  


Step 1 :   Total accumulation (before considering opportunity cost of tax cost of rolling over)  =  $46000 × 1.05620  =  $136784  

Step 2 :   ($11500) × 1.04620  = ($28270) opportunity cost for having to pay tax on rollover (by rolling over traditional IRA, Yuki will have to pay $11500 in taxes ($46,000 × 25%). Because she pays $11500 in taxes with funds that are outside her retirement accounts, she will miss the opportunity of generating a 4.6% after-tax rate of return on the $11500 for 20 years.

Step 3 :   Accumulation of Roth IRA net of opportunity cost associated with taxes paid on rollover = $108514 ($136784 - $28270).

Accumulation if she rolls traditional IRA into a Roth IRA :  $19604

Roth IRA accumulation $108514 minus traditional IRA accumulation $88910 = $19604 greater accumulation if she rolls over traditional into Roth IRA.

b.

Accumulation if she keeps funds in traditional IRA  :  $109427

109427 [$46,000 × 1.05620 × (1 - .20)]

Accumulation if she rolls funds into Roth IRA:     $108514


Step 1 :   Total accumulation (before considering opportunity cost of tax cost of rolling over)  =  $46000 × 1.05620  =  $136784  

Step 2 :   ($11500) × 1.04620 = ($28270) opportunity cost for having to pay tax on rollover (by rolling over traditional IRA, Yuki will have to pay $11500 in taxes ($46,000 × 25%). Because she pays $11500 in taxes with funds that are outside her retirement accounts, she will miss the opportunity of generating a 4.6% after-tax rate of return on the $11500 for 20 years.

Step 3 :   Accumulation of Roth IRA net of opportunity cost associated with taxes paid on rollover = $108514 ($136784 - $28270).

Accumulation if she rolls traditional IRA into a Roth IRA :  $913

When she withdraws the retirement funds in 20 years, she expects her marginal tax rate to be 20%.

$913 lower accumulation if she rolls over traditional IRA into Roth IRA (greater accumulation if she does not rollover).  See calculations below:

Roth IRA accumulation $108514 minus traditional IRA accumulation $109427 = ($913) Lower accumulation if she rolls over traditional into Roth IRA.